A chemist has one solution containing a 16% concentration of acid and a second solution containing a 26% concentration of acid. How many milliliters of each should be mixed in order to obtain 30 milliliters of a solution containing an 18% concentration of acid?

Answer by checkley77(12844)

.26x+.16(30-x)=.18*30
.26x+4.8-.16x=5.4
.10x=5.4-4.8
.10x=.6
x=.6/.10
x=6 ml. of the 26% solution is used.
30-6=24 ml. of the 16% solution is used.
Proof:
.26*6+.16*24=5.4
1.56+3.84=5.4
5.4=5.4